Question 39 1 pts A sample has M = 70 and s = 4. In this sample, what is the X value corresponding to z = X =72 O X-64 Question 40 1 pts
Answer #1

Z-score is given by

Z = X - M / S , Where M is mean and S is standard deviation.

- 1 = X - 70 / 4

X - 70 = -4

X = -4+70

X = 66